### v1

@@ 4 Discussion
In summary, we have shown here that TF embedding methods can rival state-of-the-art GNNs on the task of PSE prediction, even when trained with less data. Given the nature of the challenge, scalability should be a primary concern for anyone developing such models, and methods with linear complexity provide a huge advantage in this regard. They also bring the benefit of lower electricity consumption and therefore a reduced carbon footprint. This is, of course, an important consideration in the context of climate change. Zitnik et al. reported very poor performance of their only tested TF models, with both RESCAL and DEDICOM being outperformed even by a simple predictor. This may well explain the corresponding lack of application by later researchers. However, when set up with optimal hyperparameters and run with modern loss functions and sampling strategies such as those implemented in LibKGE, we have demonstrated that they do compete with the state-of-the-art. If the hurdle of inductive reasoning can be cleared, TF methods could be the most cost-effective way to tackle PSE prediction.
@@ 6 Author contributions statement
Oliver Lloyd: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Visualization.

### v2

@@ 4 Discussion
In summary, we have shown here that TF embedding methods can rival state-of-the-art GNNs on the task of PSE prediction, even when trained with less data. Further, we have shown that the SimplE model can be trained to almost-full power at lightning speeds. Given the nature of the challenge, scalability should be a primary concern for anyone implementing such solutions, and TF models like SimplE provide a huge advantage in this regard. They also bring the benefit of lower electricity consumption and therefore a reduced carbon footprint, which is, of course, an important consideration in the context of climate change. Zitnik et al. reported very poor performance of their only tested TF models, with both RESCAL and DEDICOM being outperformed even by a simple predictor. This may well explain the relative lack of interest in TF by later researchers. However, when set up with optimal hyperparameters and run with modern loss functions and sampling strategies such as those implemented in LibKGE, we have demonstrated that they could be more useful than the current state of the art.
@@ 6 Author contributions statement
Oliver Lloyd: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization.